OpenStudy (rock_mit182):

anyone good at calculus, Please give me a hand \[\int\limits_{}^{} e ^{xsec(e(x))^3} dx\]

OpenStudy (rock_mit182):

\[\int\limits \frac{ (1-x^2)^\frac{ 2 }{ 3 } }{ x^6 } dx\]

OpenStudy (rock_mit182):

\[\int\limits e^{xsec(e^x)^3} dx\]

OpenStudy (rock_mit182):

please someone help me out with those, two integrals

OpenStudy (anonymous):

What chapter is this?? U substitution?

OpenStudy (rock_mit182):

integral technics

OpenStudy (anonymous):

Have you learned about u sub yet?

OpenStudy (rock_mit182):

yeah

OpenStudy (rock_mit182):

U sub, by parts, partial fractions

OpenStudy (rock_mit182):

could you help me out ?

OpenStudy (rock_mit182):

IDk how to start,

OpenStudy (freckles):

@dan815 I think these integrals are too advanced for me. And too advanced for wolfram. You want to take a look?

OpenStudy (freckles):

Well at least one is too hard for wolfram. But both cannot be expressed as elementary functions.

OpenStudy (rock_mit182):

damn those integrals are trick, indeed..

OpenStudy (freckles):

This is the exact way the problems look?

OpenStudy (rock_mit182):

I think Im going to fail my midleterm

OpenStudy (anonymous):

i believe it is integration by parts, that or you would have to do a fraction decompisition...Its been a while since calc two so i cant help you out there. sorry

OpenStudy (rock_mit182):

:/ ok ...

OpenStudy (dan815):

oh ya sec is never good on e

OpenStudy (anonymous):

what book are you using and what problem is it. Sometime you can find solutions

OpenStudy (dan815):

the best u can do is expand sec as a power series and rewrite a product log definition of the integral

OpenStudy (freckles):

I wonder if the second one is actually suppose to be : \[\int\limits_{}^{}e^x \sec^3(e^x) dx \text{ if so that is tons doable }\]

OpenStudy (dan815):

otherwise u have to resort to nonelementary functions

OpenStudy (rock_mit182):

yeah i believe so is sec ^3

OpenStudy (freckles):

and the sec is not part of the exponent right?

OpenStudy (rock_mit182):

let me take a look

OpenStudy (rock_mit182):

oh yeah lucky for me it is not

OpenStudy (freckles):

\[\int\limits_{}^{}e^x \sec^3(e^x) dx \\ \text{ sub: } u=e^x \\ du=e^x dx \\ \text{ then apply integration by parts } \]

OpenStudy (rock_mit182):

sounds fair

OpenStudy (rock_mit182):

what about the first one ?

OpenStudy (freckles):

Will you confirm that is actually: \[\int\limits_{}^{}\frac{(1-x^2)^\frac{2}{3}}{x^6} dx\]

OpenStudy (rock_mit182):

100 sure

OpenStudy (anonymous):

its a partial fraction decompisistion according to wolfram http://www.wolframalpha.com/input/?i=integral+%28%281-x%5E2%29%5E%283%2F2%29%29%2Fx%5E6

OpenStudy (freckles):

if only it was 3/2 :(

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

my bad

OpenStudy (freckles):

I cannot do this because I'm not familiar with hypergeometric function . lol.

OpenStudy (anonymous):

do me a favor and punch your teacher in the face

OpenStudy (rock_mit182):